UT May 08 2022 - Split Night - DEC / PEPSI-PETS - "Night of the F/15s"



Night Info

  • Starting Configuration: LBTI-Bino
  • Sunset 19:06
  • Nautical Twilight Ends 20:04
  • Astro Twilight Ends 20:38
  • Moon sets 01:19 - Illumination 40%
  • Astro Twilight Starts 03:54
  • Nautical Twilight Starts 04:27
  • Sunrise 05:26

Staff

  • OSA: J.Williams
  • ISp; E.Solheid
  • Mtn Manager: J.Prothro
  • ISA: B.Rothberg, R.T.Gatto
  • Special Guests: LBTI

Night Plan

  • LBTI Basic Checkout - 4hrs - twilight until-00:55
    • DX checkouts (not closing the loop)
      • Pupil alignment DX
      • M3 position check DX
      • Verify bayside stages on DX
    • SX WFS software checkout (in parallel)
  • Contingency in case LBTI cannot execute Basic Checkout in good conditions
    • Configure for Bino-MODS
      • D-time science
        • SpecStd - Feige 34 - 0.5hr
        • midz-ULIRG FF0937 - 1.5hr
      • 1 hour gap
    • PEPSI Engineering continued - flux standard visible all night - 1hr
      • HR5191 RA 13hr Dec +49
  • Reconfigure to PEPSI 00:55 (23:55 - if no LBTI work done)
  • PEPSI-PETS 01:25-03:24
  • Reconfigure from PEPSI to MODS (regardless of conditions) 03:25-03:35
  • MODS D-time science ULIRG, then SpecStd

Poor Weather Backup (includes if LBTI cannot use time on-sky)

  • LUCI-Bino
    • Rothberg Local Mergers
      • OT: /home/lbto/Service_Observing/2022A_D_E/PoorWeather/Rothberg/LocalMergers/COBandhead_Mergers_V47.xml
        • 1 hour per target - 1hr15 with telluric
        • Not in order of priority - observe best placed target
        • DONE NGC 4004 - DEC +27 - twilght-02hr
        • UGC 9829 - DEC -01 - 21:30-twilight
        • NGC 3597 - DEC -23 - twilight-22hr
    • Photometric Standards - any seeing, but must be clear
      • /home/lbto/Service_Observing/2022A_D_E/PoorWeather/LUCI_Phot_Standards
      • Select stars based on visibility, need to get stars at two different airmasses
  • LBCs
    • Photometric Standards - any seeing, but must be clear
      • OT File: /home/lbto/Service_Observing/2022A_D_E/LBC/New_PhotFields/New_Phot_Fields_Master_OT_V1.1.xml
  • MODS
    • Spetrophotometric Standards
      • Observe all modes/configurations - any seeing, but must be clear

Night Log

Closed Dome

  • UT 00:48 - starting up LUCIs
  • UT 00:53 - init OK - taking test darks
    • L1 0001-0002, L2 0001-0002 - 2.51sec dark
    • putting blind masks away to save time on-sky
  • UT 00:58 - checking fieldstops on both LUCIs N375
  • UT 01:04 - done - L1 0003-0005, L2 003-0005 - blind filters in place
  • UT 01:23 - waking MODS (in case we use at end of night) on OBS2 (running two computers as backup)
  • UT 01:27 - running simSnap
  • UT 01:34 - simSnap done - OK
    • MODS and LUCI ready for tonight (no plans to use LBCs)
  • UT 01:35 - winds over limits in the last 20min - thick clouds have thinned out
  • UT 01:57 - polycom issues - first RM507 connected to summit, but BRothberg couldn't. Then J.Williams rebooted it, which allowed B.Rothberg to connect, but RM507 could not connect. RM507 now rebooting polycom.
  • UT 02:01 - reboot of all polycoms seems to have worked - clouds thinned out - but winds reaching 26m/s
  • UT 02:15 - Telescope remaind closed at sunset due to high winds
  • UT 03:07 - telescope still closed due to winds over limits - any remaining time before PEPSI/PETS will be given to LBTI
  • UT 04:03 - mods1data alert - IT 7285 - looks like this is the same issue - mods were awake, i had taken simSnaps but otherwise not using it
    • Trigger time: 04:03:37
  • UT 04:19 - restarted services as per IT - taking test biases on MODS1
  • Proceeded with restarting services as noted in this IT - taking test bias exposures.
    • M1B reads out fine, M1R doing the slow readout - and need to do a red fitsflush
    • - could just be coincidence?
    • M1R had two bias images with delayed reading out
    • Crap... forgot the three digit extension fix
    • isisCmd --mods1 m1.rc use mount 3
      STATUS: USE Path=UNIX:\CB\/lhome/data
      hit CTRL C to return to prompt
    • OK...now back to our regularly scheduled 4 digits for filenumbering
  • UT 04:35 - winds trending downwards it seems

ON-Sky

LBTI Tests
  • UT 04:41 - opening dome for LBTI SX/DX Eng (will run until winds/weather or PEPSI/PETS handover)
  • UT 05:18 - A.Vaz noted that the star on DX seems visibly brighter than SX - will try and get measured mags (if not tonight then later this week)
  • Comparison_SX_vs_DX_LBTI.png
  • UT 05:59 - tests seemed to go successfully for LBTI - may hand over early - will use LUCI for filler science before PEPSI/PETS
Bino-LUCIs filler science
  • UT 23:07 - reconfiguring for bino-LUCIs - will attempt some science before handover to PEPSI/PETS
  • UT 06:10 - pointing check
  • UT 06:15 - collimation preset
NGC 4004
  • UT 06:18 - sending telluric preset
  • Looks like LUCI-1 is out of focus - star is a blob and longslit is out of focus
  • L2 -9.66", -2.04"
    • Move to N1.8 and back to N375 to check focus on LUC1
    • forgot to put clear+J in to see slit
    • this worked, slit is in focus as is the star throught the slit
  • L1 -10.2", -1.67"
    • RTD Y axis is in reverse order - moving cursor down it goes to higher Y values
    • PA is 180 - could this be why?
    • send terminal dump to G.Eychaner
  • UT 06:28 - starting spectroscopy
    • L1 0014-0015
    • L2 0009-0010
  • Stopped using test rtds
  • UT 06:31 - particle counts hitting limts
  • UT 06:32 - sending preset to NGC 4004
  • UT 06:36 - had error inst commit for L1 but not clear what it was
    • Another detector focus error
    • 2022.05.08 06:32:22 s | error medium | luci.luc | Luci ONE | ONE_InstrumentM | RMIInstrumentManagerImpl.java#commitSetu(323) | error while committing setup DETECTOR_FOCUS
  • UT 06:38 - closing up - putting LUCIs in safe config - particle counts too high
  • Updated IT 8614 about detector focus.
  • UT 07:00 - counts down - opening up again
  • UT 07:04 - moving the detector focus 700 steps until limit switch comes on
    • at neg limit now
  • UT 07:07 - sending preset
    • L1 -0.3428", -2.1852"
    • L2 1.9218", -2.55"
  • UT 07:21 - starting spectroscopy
    • L2 GEIRS error - doing manually
    • 2022.05.08 07:21:50 s | user medium | luci.luc | Luci ONE | ONE_HIRAMO | RMIHIRAMOImpl.java#activateLoop(895) | Closed loop operation activated for HIRAMO card at 192.168.0.10(192.168.0.10):4009.
      2022.05.08 07:22:11 s | error medium | luci.luc | Luci TWO | TWO_ReadoutMana | GEIRSClient.java#setROEMode(188) | Could not commit value "MER, 10" by "setROEMode" to GEIRS from "rmi://localhost:60002/TWO_GEIRSServer"!
      2022.05.08 07:22:12 s | error high | luci.luc | Luci ONE | ONE_GEIRSServer | RMIGEIRSServiceImpl.java#checkOkAnswer(946) | GEIRS received response with error: ERROR 19 Command 'subwin off' returned errorcode = 19: (E_invcamera=19) invalid ROE response: 'ERROR in extend_camera_cmd (/home/readout1/GEIRS/trunk-r770M-15/pttrns/Luci1/roe_rdmode_mer.luci1:8)'
    • Updating IT8603
  • SX guider 1.4", DX guider is 1.14"
  • UT 07:42 - Particle counts are borderline
  • UT 07:44 - closing for winds - aborted mid exposure
    • L1 0023-0025
    • L2 0017-0019
  • UT 07:45 - putting LUCIs in safe config

PEPSI-PETS

  • Target is WASP 103 RA 16hr Dec +07 - this will likely put us into wind once it transits
  • UT 07:53 - reconfiguring for PEPSI
  • UT 08:50 - opening dome
  • UT 09:03 - can't see any star on GCS - fiber movement by I.ilyin - still nothing
  • UT 09:07 - moved to brighter star - slightly better - pointing was way off
  • UT 10:24 - got 1hr on target - facing into wind, but haven't hit limits
  • SX vs DX flux comparison on blue
  • PEPSI_Comparison_Flux_SX_vs_DX.png

D-time Science

IR17044+6720
  • UT 10:28 - reconfigure for bino-MODS - will go for as much as possible then specstd
  • the LUCI target was at Dec -01 which would have been worser than the Dec +7 for PETS
  • UT 10:41 - pointing preset
  • UT 10:45 - collimation preset
  • UT 10:47 - sending preset to science target
    • guiders show 1"-1.1"
    • MODS2 Offset Command:
    • offsetxy 4.245 8.595 rel
    • MODS1 Offset Command:
    • offsetxy 0.080 11.968 rel
    • offsetting M1 -0.123" DET X
  • UT 10:59 - staring science
  • UT 11:02 - Guiders show 1.1"
  • UT 11:34 - sending preset to BD+33
    • MODS-1 looks stuck on preset
    • MODS2 Offset Command:
    • offsetxy 4.071 8.927 rel
    • Resending pereset on MODS-1 - got stuck on set object
    • MODS-2 tweak
    • MODS2 Offset Command:
    • offsetxy -0.466 -0.287 rel
  • UT 11:42 - starting MODS-2 spectroscopy
    • M2B 0004-0005
    • M2R 0009-0010
    • MODS1 Offset Command:
    • offsetxy -0.979 12.245 rel
  • UT 11:45 - starting MODS-1 spectroscopy
  • UT 11:54 - got those last photons!
  • UT 11:56 - closing up

Calibrations

  • Will use AZ partner MODS cals (Dual Grating slitless flats and arcs)
    • UT May 04
      • M1B 0013-0015 - Arcs
      • M1R 0013-0015 - Arcs
      • M2B 0013-0015 - Arcs
      • M2R 0013-0015 - Arcs
    • UT May 06
      • M1B 0019-0028 - slitless flats
      • M1R 0016-0020 - slitless flats
      • M2B 0019-0028 - slitless flats
      • M2R 0016-0020 - slitless flats
  • UT 12:04 - putting both MODS to sleep
  • LUCI Cals G210 2.27microns 0.5" slit Kband
    • L1 0026-0034 - do not use
      • script seems too bright for flats - saturating?
      • yes, the DIT was set to 40 in fraternal library, but for L1 its 20sec
      • also different halos, halo1 was used with 40sec, in mono, its halo 2 with 20sec
      • fixed and restarted - counts are 5k per integrated frame - so running this part of script 2x rather than futz with halo1
    • L1 0035-0054 - flats 2x
    • L1 0055-0062 - arcs
    • L2 0020-0037
  • UT 12:24 - putting LUCI-2 into safe config
  • UT 13:02 - putting LUCI-1 into safe config

-- BarryRothberg - 03 May 2022
Topic revision: r13 - 08 May 2022, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback